OpenStudy (anonymous):

PLEASE HELP ME If 2 lines have the same slope but different y-intercepts, then they will have __ solutions a. none b. one c. infinitely many d. cannot be determined Monica and Michael both want to buy a scooter. Monica has already saved $25 and plans to save $5 per week. Michael has $16 and plans to save $8 per week. In how many weeks will they have the same amount of money saved? a. one b. two c. three d. four Find all solutions for 2y - 4x = 2 and y = 2x + 1. a. none b. infinitely many c. (0, 4) d. (2, 5)

OpenStudy (anonymous):

@sword

OpenStudy (kearsten):

the answer to the first question is a i think

OpenStudy (anonymous):

ok thnaks do u know the other

OpenStudy (anonymous):

thanks

OpenStudy (kearsten):

im not very good at math i only knew the first sry

OpenStudy (anonymous):

thanks anyway

OpenStudy (anonymous):

@flashp

OpenStudy (anonymous):

@nickmeneses1997

OpenStudy (anonymous):

@CountryGurl15

OpenStudy (loser66):

for a) you are correct, 2 parallel lines has no solution

OpenStudy (loser66):

for b) the equation is 25+5x= 16+8x . solve for x to get the answer

OpenStudy (loser66):

for c) isolate y from the first one , it is y = 2x+1, hence it is exactly the same with the second one. So, the answer is infinitely many solution.

OpenStudy (anonymous):

for b) 3 weeks for c)y=0.5x+1=2x+1 1.5x=0 so x=0 put in either equation and y=1 answer(0,1) intersection